1. Suppose that computers use 4 units of capital for each worker, so that KC = 4 ⋅ LC , whereas shoes use 0.2 units of capital for each worker, so that KS = 0.2 ⋅ LS. There are 200 workers and 200 units of capital in the economy.

a.Solve for the amount of labor and capital used in each industry.

b. Suppose that the number of workers increases to 250 because of immigration, keeping total

capital fixed at 200. Again, solve for the amount of labor and capital used in each industry.

c. Suppose instead that the amount of capital increases to 250 due to FDI, keeping the total

number of workers fixed at 200. Again solve for the amount of labor and capital used in

each industry.

d. Explain how your results in parts (b) and (c) are related to the Rybczynski theorem.
Answer #1

a) The box diagram shows that the amount of labour and capital used in each industry must add up to the total for the economy,so that Kc+Ks=200 and Lc+Ls=200

Use the facts that Kc=4 x Lc and Ks=0.2xLs to rewrite these equations are:

4 x Lc+0.2xLs= 200 and Lc+Ls=200

Use these two equations to solve for Lc and Ls and then calculate the amount of capital used in each industry using Kc =4 x Lc and Ks = 250

c)Kc. = 4 * Lc and Ks = 0.2 * Ls workers or labours are 200 whereas there are now 250 units of capital.so Kc + Ks = 250 and Lc +Ls=200 substitue for Kc

d)this can leads to adjustment in the quantities and types of outputs between the two regions.it explains the outcome from an increase in one of these factor's supply as well as the effect on the out put of a good which depends on an oppositing factor.